# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ 3.2 Weight calibration (SIMCal-W: unit-mean ratios via OOF stacking of $S$-monotone candidates)
    Monotone projections (train on $I_{\neg k}$). Fit increasing and decreasing isotonic maps on $S$ (the latter by running PAVA on S-S), rescale each to mean one on $I_{\neg k}$, and predict OOF candidates on $I_{k}$: $W_{\uparrow}^{\mathrm{OOF}}$, $W_{\downarrow}^{\mathrm{OOF}}$; optionally include an identity candidate $W_{\mathrm{base}}^{\mathrm{OOF}}\equiv 1$.
@@
    OOF stacking (variance-aware). Define residuals $\Delta_{i}$ used by the downstream estimator: $\Delta_{i}{=}R_{i}$ for IPS and $\Delta_{i}{=}R_{i}-\hat{m}(X_{i},A_{i})$ for DR (with $\hat{m}{=}\hat{q}$ below). Let $U_{c}=W_{c}^{\mathrm{OOF}}\Delta$ for $c\in\{\mathrm{base},\uparrow,\downarrow\}$ and compute $\hat{\Sigma}_{cd}=\mathrm{cov}(U_{c},U_{d})$ (tiny ridge if needed). Choose simplex weights
@@
    Light variance guard (optional; $\rho{=}1$ by default). Cap variance at absolute threshold $\rho$:
@@
Remark (transport view). In the continuous case the ideal component is $m^{\star}(s)=p_{S\mid\pi^{\prime}}(s)/p_{S\mid\pi_{0}}(s)$; SIMCal-W stacks increasing and decreasing $S$-monotone candidates to minimize IF variance, yielding approximately monotone weights when one direction dominates. When $m^{\star}(S)$ is not approximately monotone, isotonic projection trades bias for variance; in the low-overlap regimes where this matters most, we recommend Direct over IPS/DR (see CLE diagnostic, Section 2).
[…]
3.3 Estimators: Calibrated IPS and DR-CPO
@@
DR-CPO (Doubly Robust Calibrated Policy Outcomes).
@@
3.4 IF-space stacking (variance-optimal convex ensembling)
@@
3.5 Oracle-Uncertainty-Aware (OUA) inference
Standard CIs treat the calibration function $\hat{f}$ as fixed, ignoring first-stage estimation error. When the oracle slice is small, this understates total variance and produces dramatically narrow intervals. In our experiments, naive CIs on uncalibrated scores achieve 0% coverage; bootstrap inference with bias-corrected estimation achieves $\sim$}95% coverage for Direct mode.
OUA variance decomposition.
@@ 3.5 Oracle-Uncertainty-Aware (OUA) inference
OUA confidence intervals.
@@
The OUA share $\widehat{\operatorname{Var}}_{\text{cal}}/\widehat{\operatorname{Var}}_{\text{total}}$ ranges from 0% (100% oracle fraction, where no calibration is needed) to over 50% (5% oracle fraction), explaining why ignoring it produces catastrophically narrow CIs.
@@
While the jackknife captures variance, it does not correct for bias in the plug-in estimator $\widehat{V}=n^{-1}\sum_{i}\hat{f}(S_{i})$. This bias arises because $\hat{f}$ is fit on the oracle slice and applied to all samples, creating a covariance between the calibration function and the evaluation samples. We address this with the one-step estimator derived from the efficient influence function for missing-outcome estimation (Section C.13):
@@
where $\hat{f}$ is the full calibrator, $\hat{f}^{(-i)}$ is the calibrator trained on folds excluding sample $i$’s cluster, and $L$ is the oracle slice. The first term uses the full calibrator for lower variance; the second term uses cluster-out-of-fold predictions to avoid overfitting in the residual correction. If oracle sampling is non-uniform (e.g., oversampling uncertain prompts), replace the constant $|L|/n$ with sample-specific inclusion probabilities $e(Z_{i})$ in the general AIPW form (Section C.13). Combined with bootstrap (refitting the calibrator in each replicate), this achieves $\sim$}95% coverage across oracle fractions (Table 3), compared to 70–89% for jackknife-only OUA.
@@
Bootstrap with $\hat{\theta}_{\mathrm{aug}}$ requires $B$ refits of AutoCal-R (we use B=2, [^0] B{=}2{,}000). Since isotonic regression is $O(n\log n)$, the overhead remains modest relative to the cost of teacher forcing and response generation. The full bootstrap procedure (Algorithm 7) uses prompt-level cluster resampling to preserve within-prompt dependence (Cameron et al., 2008), with a resample-until-valid loop ensuring each replicate has sufficient oracle labels.
4 Theory: EIF, Design-by-Projection, and Efficiency
@@
Let $L^{2}_{0}$ be the mean-zero Hilbert space with inner product $\langle f,g\rangle=\mathbb{E}[fg]$. Let $\mathcal{M}$ be a baseline semiparametric model with tangent space $T(P)$ and canonical gradient (EIF) $\phi^{\star}$. When justified knowledge defines a restricted model $\mathcal{M}_{\mathcal{C}}\subset\mathcal{M}$ (e.g., mean sufficiency in $S$, monotonicity of conditional expectations), the restricted tangent space $T_{\mathcal{C}}(P)\subseteq T(P)$ yields a new canonical gradient $\phi^{\star}_{\mathcal{C}}$.
@@
Remark (prior art). Theorem 2 formalizes a well-known principle in semiparametric efficiency theory (Newey, 1990; Bickel et al., 1993); our contribution is naming it and systematically applying it to surrogate-based LLM evaluation via the Design-by-Projection framework.
Corollary 1 (Blackwell–efficiency monotonicity).
If $S_{2}$ is a garbling of $S_{1}$ (i.e., $\sigma(S_{2})\subseteq\sigma(S_{1})$), then $\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{1})\big)\leq\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{2})\big)$, with strict inequality unless $W_{\pi^{\prime}}$ is already $\sigma(S_{2})$-measurable and $R^{\star}$ is degenerate.
@@
(i) Conditioning: taking $\mathcal{C}=\{f:\,f=\mathbb{E}[f\mid S]\}$ recovers Theorem 1. (ii) Mean-one monotone weights: restricting the weight component to $\{w:\mathbb{E}[w]=1,\;w\uparrow S\}$ motivates SIMCal-W; the exact $\mathrm{IsoMeanOne}_{S}$ projection weakly reduces dispersion in finite samples by majorization. (iii) Stacking: restricting to the convex hull of candidate IF columns gives the variance-optimal convex ensemble.
@@
Let $R=\hat{f}(Z(S))$ be AutoCal-R (monotone in $S$ or two-stage index; cross-fitted), and let $W^{\mathrm{m1}}_{\pi^{\prime}}\!\triangleq W_{\pi^{\prime}}/\bar{W}_{n}$ denote the self-normalized (SNIPS) ratios, where $\bar{W}_{n}=n^{-1}\sum_{i}W_{\pi^{\prime},i}$. Let $\hat{W}_{\pi^{\prime}}$ be SIMCal-W weights (OOF stack ++ variance guard $\rho\geq 1$). Then $\widehat{V}_{\mathrm{IPS}}=\tfrac{1}{n}\sum_{i}\hat{W}_{\pi^{\prime},i}R_{i}\to_{p}V(\pi^{\prime})$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection (see App. A.3), $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ and $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$; the reference implementation uses a simpler normalization that empirically yields similar ESS gains.
Theorem 3 (DR-CPO: $\sqrt{n}$ limits and efficiency).
@@
i.e., DR-CPO attains the surrogate efficiency bound. (Efficiency attainment under the product-rate condition is a standard property of DR estimators (Bang & Robins, 2005; Chernozhukov et al., 2018); our contribution is verifying that the CJE construction satisfies these conditions.)
Budgeted bound (variance cap).
For $\rho\geq 1$, define
Theorem 4 (Budgeted information bound).
The optimal asymptotic variance under the cap $\rho$ equals $\operatorname{Var}(\phi^{(\rho)})$, which is nonincreasing in $\rho$ and satisfies $\lim_{\rho\to\infty}\operatorname{Var}(\phi^{(\rho)})=\operatorname{Var}(\phi_{\mathrm{sur}})$. If SIMCal-W converges to $m^{\star}_{\rho}$ (with the guard), then DR-CPO attains $\operatorname{Var}(\phi^{(\rho)})$.
Theorem 5 (IF-space stacking).
Corollary 2 (Carathéodory sparsity).
Proposition 2 (OUA jackknife variance estimation).
@@
Data requirement. Applying this test requires a small oracle slice under each target policy $\pi^{\prime}$ (or per evaluation environment, e.g., time window or subgroup). We recommend collecting a modest audit sample (${\sim}50$200200 labels per policy) to validate that the base-trained calibrator can be safely reused; if a policy fails the test, either recalibrate with policy-specific data or fall back to oracle-only evaluation for that cell.
@@ Discussion.
Theorem 2 formalizes Design-by-Projection: when justified knowledge defines a restricted statistical model (smaller tangent space), the efficiency bound in the restricted model is at most the bound in the baseline model. The key instantiations are: Theorem 1 (conditioning on $S$), Theorem 4 (budgeted monotone weights), and Theorem 5 (convex hull of candidate IFs). AutoCal-R, SIMCal-W, and DR/TMLE with cross-fitting attain the corresponding bounds in their respective restricted models. In finite samples, the exact $\mathrm{IsoMeanOne}_{S}$ projection additionally majorizes dispersion (Lemma 2), explaining the ESS gains delivered by SIMCal-W.
@@
(i) Mean sufficiency generalization. Theorems 1 and 3 assume mean sufficiency in $S$: $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$. AutoCal-R’s two-stage mode uses a learned index $Z(S,X)$; the theory generalizes by replacing $S$ with $Z(S,X)$ throughout, provided $Z$ is consistently estimated. (ii) Double robustness and weight misspecification. Theorem 3 assumes at least one nuisance converges at rate $n^{-1/4}$. When teacher-forced weights are unreliable (as in high-overlap-failure regimes), DR remains consistent via the outcome model alone; however, the efficiency bound $\operatorname{Var}(\phi_{\mathrm{sur}})$ is attained only when both nuisances are well-specified. In practice, when IPS fails due to poor coverage, DR effectively reduces to the Direct Method, which achieves the same consistency without the variance penalty from unstable weights (see Section 5, Q3). (iii) SIMCal-W convergence. Theorem 4 assumes SIMCal-W converges to $m^{\star}_{\rho}$; this holds under standard regularity conditions for isotonic regression and cross-validated stacking (see, e.g., van der Laan et al. (2007)).

### v3

@@ 3.2 Weight stabilization (unit-mean ratios via OOF stacking of $S$-monotone candidates)
    Monotone projections (train on $I_{\neg k}$): Fit increasing and decreasing isotonic maps on $S$ (the latter by running PAVA on S-S), rescale each to mean one on $I_{\neg k}$, and predict OOF candidates on $I_{k}$: $W_{\uparrow}^{\mathrm{OOF}}$, $W_{\downarrow}^{\mathrm{OOF}}$; optionally include an identity candidate $W_{\mathrm{base}}^{\mathrm{OOF}}\equiv 1$.
@@
    OOF stacking (variance-aware): Define residuals $\Delta_{i}$ used by the downstream estimator: $\Delta_{i}{=}R_{i}$ for IPS and $\Delta_{i}{=}R_{i}-\hat{m}(X_{i},A_{i})$ for DR (with $\hat{m}{=}\hat{q}$ below). Let $U_{c}=W_{c}^{\mathrm{OOF}}\Delta$ for $c\in\{\mathrm{base},\uparrow,\downarrow\}$ and compute $\hat{\Sigma}_{cd}=\mathrm{cov}(U_{c},U_{d})$ (tiny ridge if needed). Choose simplex weights
@@
    Light variance guard (optional; $\rho{=}1$ by default): Cap variance at absolute threshold $\rho$:
@@
Remark (transport view). In the continuous case the ideal component is $m^{\star}(s)=p_{S\mid\pi^{\prime}}(s)/p_{S\mid\pi_{0}}(s)$; weight stabilization stacks increasing and decreasing $S$-monotone candidates to minimize IF variance, yielding approximately monotone weights when one direction dominates. When $m^{\star}(S)$ is not approximately monotone, isotonic projection trades bias for variance; in the low-overlap regimes where this matters most, we recommend Direct over IPS/DR (see CLE bound, Section 2). (Weight stabilization visualization in Section D.9.)
3.3 Estimators: Direct, Calibrated IPS, and Calibrated DR
@@
Direct (fresh responses under $\pi^{\prime}$; no overlap needed).
When responses can be generated under the target policy, the Direct estimator uses calibrated rewards without importance weighting:
$\widehat{V}_{\mathrm{Direct}}(\pi^{\prime})=\frac{1}{n}\sum_{i=1}^{n}R_{i}=\frac{1}{n}\sum_{i=1}^{n}\hat{f}(Z_{i}).$
For inference, the bias-corrected augmented estimator $\hat{\theta}_{\mathrm{aug}}$ (Equation 3) achieves $\sim$}95% coverage by correcting for calibration uncertainty. Direct requires no overlap assumption and is the recommended default when fresh generation is feasible (Table 7).
@@
Calibrated DR (doubly robust with calibrated rewards).
@@
Remark (DR vs Direct). DR combines an outcome model with a weighted residual correction. Under good overlap, the IPS component can improve efficiency; under poor overlap (TTC <0.7<0.7), it adds noise. In our experiments, DR effectively reduces to its outcome model, and Direct achieves the same accuracy without requiring logprobs (Section 5, Q3).
3.4 Influence-function stacking (variance-optimal convex ensembling)
@@
3.5 Calibration-aware inference
Standard CIs treat the calibration function $\hat{f}$ as fixed, ignoring first-stage estimation error. When the oracle slice is small, this understates total variance and produces severely narrow intervals. In our experiments, naive CIs on uncalibrated scores achieve 0% coverage; bootstrap inference with bias-corrected estimation achieves $\sim$}95% coverage for Direct mode.
Variance decomposition.
@@ 3.5 Calibration-aware inference
Calibration-aware confidence intervals.
@@
The calibration uncertainty share $\widehat{\operatorname{Var}}_{\text{cal}}/\widehat{\operatorname{Var}}_{\text{total}}$ ranges from 0% (100% oracle fraction, where no calibration is needed) to over 50% (5% oracle fraction), explaining why ignoring it produces catastrophically narrow CIs.
@@
While the jackknife captures variance, it does not correct for bias in the plug-in estimator $\widehat{V}=n^{-1}\sum_{i}\hat{f}(S_{i})$. This bias arises because $\hat{f}$ is fit on the oracle slice and applied to all samples, creating a covariance between the calibration function and the evaluation samples. We address this with the one-step estimator derived from the efficient influence function for missing-outcome estimation (Section C.14; see also Chen et al. (2026) for the LLM-as-judge instantiation):
@@
where $\hat{f}$ is the full calibrator, $\hat{f}^{(-i)}$ is the calibrator trained on folds excluding sample $i$’s cluster, and $L$ is the oracle slice. The first term uses the full calibrator for lower variance; the second term uses cluster-out-of-fold predictions to avoid overfitting in the residual correction. If oracle sampling is non-uniform (e.g., oversampling uncertain prompts), replace the constant $|L|/n$ with sample-specific inclusion probabilities $e(Z_{i})$ in the general AIPW form (Section C.14). Combined with bootstrap (refitting the calibrator in each replicate), this achieves $\sim$}95% coverage across oracle fractions (Table 3), compared to 70–89% for jackknife-only variance estimation.
@@
Bootstrap with $\hat{\theta}_{\mathrm{aug}}$ requires $B$ refits of the reward calibrator (we use B=2, [^0] B{=}2{,}000). Since isotonic regression is $O(n\log n)$, the overhead remains modest relative to the cost of teacher forcing and response generation. The full bootstrap procedure (Algorithm 7) uses prompt-level cluster resampling to preserve within-prompt dependence (Cameron et al., 2008), with a resample-until-valid loop ensuring each replicate has sufficient oracle labels.
4 Theory: EIF, Model Restriction, and Efficiency
@@
We now formalize a standard principle from semiparametric efficiency theory and show how it unifies the CJE projections. Let $L^{2}_{0}$ be the mean-zero Hilbert space with inner product $\langle f,g\rangle=\mathbb{E}[fg]$. Let $\mathcal{M}$ be a baseline semiparametric model with tangent space $T(P)$ and canonical gradient (EIF) $\phi^{\star}$. When justified knowledge defines a restricted model $\mathcal{M}_{\mathcal{C}}\subset\mathcal{M}$ (e.g., mean sufficiency in $S$, monotonicity of conditional expectations), the restricted tangent space $T_{\mathcal{C}}(P)\subseteq T(P)$ yields a new canonical gradient $\phi^{\star}_{\mathcal{C}}$.
@@
Remark (prior art and novelty). The efficiency improvement from model restriction is well-known in semiparametric theory (Newey, 1990; Bickel et al., 1993). Our contribution is not the principle itself but its systematic application: we show that reward calibration, weight stabilization, and influence-function stacking are all instances of model restriction, and we provide the operational machinery (transport audit, CLE bound and TTC/AB diagnostics, calibration-aware inference) that makes the theory usable for LLM evaluation.
@@
(i) Conditioning: taking $\mathcal{C}=\{f:\,f=\mathbb{E}[f\mid S]\}$ recovers Theorem 1. (ii) Mean-one monotone weights: restricting the weight component to $\{w:\mathbb{E}[w]=1,\;w\uparrow S\}$ motivates weight stabilization; the exact $\mathrm{IsoMeanOne}_{S}$ projection weakly reduces dispersion in finite samples by majorization. (iii) Stacking: restricting to the convex hull of candidate IF columns gives the variance-optimal convex ensemble.
Direct mode: missing-data estimation.
When fresh responses can be generated under $\pi^{\prime}$, evaluation becomes a missing-data problem: we observe cheap surrogate predictions $\hat{\mu}(Z)$ for all samples but expensive oracle labels $Y$ only for a labeled subset $L$.
Proposition 1 (Direct: EIF and one-step estimator).
Let $\theta:=\mathbb{E}_{\pi^{\prime}}[Y]$ be the policy value. Under MAR ($L\perp Y\mid Z$) with $e(z):=\Pr(L{=}1\mid Z{=}z)$, the efficient influence function is
$\phi_{\mathrm{Direct}}(O)=\mu(Z)-\theta+\frac{L}{e(Z)}\big(Y-\mu(Z)\big),$
where $\mu(z)=\mathbb{E}[Y\mid Z{=}z]$. Under approximately uniform labeling ($e\approx|L|/n$), the one-step estimator is $\hat{\theta}_{\mathrm{aug}}$ (Equation 3), which attains the semiparametric efficiency bound.
In plain terms: Direct is the recommended default when fresh generation is feasible. It requires no overlap assumption and achieves optimal efficiency for missing-data estimation. (Proof in Section C.14.)
OPE mode: reweighting logged data.
When fresh generation is unavailable, importance-weighted estimators reweight logged data from $\pi_{0}$ to estimate $V(\pi^{\prime})$. The following results apply to this regime.
@@
Let $R=\hat{f}(Z(S))$ be the reward calibrator (monotone in $S$ or two-stage index; cross-fitted), and let $W^{\mathrm{m1}}_{\pi^{\prime}}\!\triangleq W_{\pi^{\prime}}/\bar{W}_{n}$ denote the self-normalized (SNIPS) ratios, where $\bar{W}_{n}=n^{-1}\sum_{i}W_{\pi^{\prime},i}$. Let $\hat{W}_{\pi^{\prime}}$ be stabilized weights (OOF stack ++ variance guard $\rho\geq 1$). Then $\widehat{V}_{\mathrm{IPS}}=\tfrac{1}{n}\sum_{i}\hat{W}_{\pi^{\prime},i}R_{i}\to_{p}V(\pi^{\prime})$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection (see App. A.3), $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ and $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$; the reference implementation uses a simpler normalization that empirically yields similar ESS gains.
Theorem 3 (Calibrated DR: $\sqrt{n}$ limits and efficiency).
@@
i.e., calibrated DR attains the surrogate efficiency bound. (Efficiency attainment under the product-rate condition is a standard property of DR estimators (Bang & Robins, 2005; Chernozhukov et al., 2018); our contribution is verifying that the CJE construction satisfies these conditions.)
Additional results (in appendix).
The budgeted information bound (Theorem 4) formalizes variance caps on importance weights: limiting weight variance ($\rho$) trades a bit of asymptotic efficiency for much better finite-sample behavior. Influence-function stacking (Theorem 5) shows that variance-optimal convex combinations of estimators achieve variance no worse than the best component. Both results apply to OPE regimes (IPS/DR); Direct mode does not use weights and is unaffected. Full statements and proofs are in Section C.18.
Proposition 3 (Calibration-aware jackknife variance estimation).
@@
Data requirement. Applying this test requires a small oracle slice under each target policy $\pi^{\prime}$ (or per evaluation environment, e.g., time window or subgroup). We recommend collecting a modest audit sample (${\sim}50$$200$ labels per policy) to validate that the base-trained calibrator can be safely reused; if a policy fails the test, either recalibrate with policy-specific data or fall back to oracle-only evaluation for that cell.
@@ Discussion.
CJE addresses two distinct statistical problems: (i) missing-data estimation (Direct mode, 1), where fresh responses are generated under $\pi^{\prime}$ and the only missing data are oracle labels; and (ii) off-policy evaluation (IPS/DR modes), where logged data from $\pi_{0}$ must be reweighted to estimate $V(\pi^{\prime})$. Direct is the recommended default when fresh generation is feasible: it requires no overlap assumption, achieves the semiparametric efficiency bound for missing-data estimation, and attains the best ranking accuracy in our experiments. For OPE regimes, Theorem 2 formalizes the model restriction principle: encoding justified knowledge (conditioning on $S$, monotone weights, convex IF combinations) tightens efficiency bounds. The budgeted bound and IF stacking theorems (Section C.18) provide additional variance control for IPS/DR. In finite samples, the exact $\mathrm{IsoMeanOne}_{S}$ projection additionally majorizes dispersion (Lemma 2), explaining the ESS gains delivered by weight stabilization.
@@
(i) Mean sufficiency and transport. The EIF derivations (Theorems 1 and 3) assume mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$), a structural assumption enabling the efficiency theory. For valid estimation, the fundamental requirement is transport ($\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$), which is strictly weaker: mean sufficiency implies transport, but transport can hold (and be verified via audit) without it. The two-stage calibration mode uses a learned index $Z(S,X)$; the theory generalizes by replacing $S$ with $Z(S,X)$ throughout, provided $Z$ is consistently estimated. (ii) Double robustness and weight misspecification. Theorem 3 assumes at least one nuisance converges at rate $n^{-1/4}$. When teacher-forced weights are unreliable (as in high-overlap-failure regimes), DR remains consistent via the outcome model alone; however, the efficiency bound $\operatorname{Var}(\phi_{\mathrm{sur}})$ is attained only when both nuisances are well-specified. In practice, when IPS fails due to poor coverage, DR effectively reduces to the Direct Method, which achieves the same consistency without the variance penalty from unstable weights (see Section 5, Q3). (iii) Weight stabilization convergence. Theorem 4 assumes weight stabilization converges to $m^{\star}_{\rho}$; this holds under standard regularity conditions for isotonic regression and cross-validated stacking (see, e.g., van der Laan et al. (2007)).